Suppose you borrowed a certain amount of money 382 weeks ago at an annual interest rate of 3% with semiannual compounding (2 times per year). If you returned $10,456 today, how much did you borrow?
Answer #1

Formula:
Present value = Future value÷(1+Interest rate)^n

PV= FV÷(1+r)^n

n = no. of periods = 382/26 = 14.662308
r = 3/2 = 1.5% per semi annum.

PV = $10,456÷(1.015)^14.662308
PV =$ 8401.65


how much did you borrow=$ 8401.65